FEL Upgrade Project Weekly Brief
April 23-27, 2001

On Saturday the FEL team and approximately half of the Jefferson Lab staff educated (and entertained) about
5,000 visitors from the local area during our 2001 Open House.  We had a continuous stream of visitors from
10:00am to 4:00pm who were shown the FEL, exhibits in three of the user labs on basic and applied applications
of the FEL, and an exhibit on the planned addition of the Helios synchrotron to FEL Facility.

On the upgrade front, we now have  successful model of our small dipoles that meet the field uniformity
requirements  (1/10,000).

The FEL program at Jefferson Lab was given high visibility at a Congressional briefing for the Virginia delegation
on April 26 which was organized by the Virginia Research and Technology Advisory Commission (VRTAC).
This is the third year that VRTAC has prioritized R&D activities in Virginia for the benefit of the delegation.  The
commission is co-chaired by Adm. Marc Pelaez, a former Chief of Naval Research and  Prof. Anita Jones
(UVA), the former assistant secretary of RDT&E for DOD.  Prior to VRTAC 's briefing on April 26, the current
CNR, Adm. Jay Cohen addressed the commission and again stressed his agency's interest in continuing FEL

On Friday, April 27, we were visited by two current members of the Defense Science Board panel on high
energy lasers (Dr. Darrell Collier and Dr. Steve Hadley for a briefing and tour of the FEL.

WBS 3 (Beam Physics):
Refinement of magnet specifications, in collaboration with WBS 9 (Biallas) and consultation with AES/DULY
continues.  Fine-tuning of the energy recovery arc telescope design to accommodate optical cavity diagnostics
and x-ray harvesting apparatus approaches completion; layouts will go to engineering "soon".  A geometric
quasi-perturbative model of the energy recovery arc is being generated to cross-check DIMAD acceptance

WBS 4 (Injector):
Continued design work on the space between the gun and lightbox. Ordered a 100mm aperture RF shielded gate
valve for isolating the gun. Details for the cathode support tube modification have been completed.

WBS 6 (RF):
Zone 3 -  Waiting on parts.  The Circulators for this zone have been received and are being tested.  They will be
installed later this summer.

Zone 4 - Complete, except for a circulator being repaired and 8 RF Control Modules were removed for the
CEBAF accelerator.

Injector Quarter - The first article 100 kW Klystron is scheduled for factory tests on June 20th.  The design
drawings of the LCW manifolds for the klystron carts were signed today.  The 100 kW Circulators have been
received.  The  vendor for the HVPS upgrade has still not decided to accept the contract.  I am making plans to
perform the HVPS upgrade in-house.

WBS 8 (I&C):
The 4-channel VME timing/delay module we are developing is proceeding well. This card is meant to satisfy all
of the upgrade general purpose timing requirements as well as being an integral part of the Drive Laser Pulse
Controller (DLPC).  It will have the following list of features: a) Four independently programmable outputs,
where each output is a two-variable function of the common event trigger input. The first variable being a delay
time after the trigger and the second variable being a duration value of the delayed event. b) The primary input
trigger will the product of a highly-configurable front-end circuit that offers both fiber (820nm) and TTL (50 ohm)
trigger inputs. c) All four outputs can be made synchronous to an externally provided master timing reference
signal (optionally used). d) All counters will be configurable to provide the best combination of max-count-time
and resolution, where max counter resolution can range from +/-50ns (5ms max delay) to +/-50ms (100 min.
max delay) on each channel independently. e) The card will be most commonly controlled by using a
standardized (within FEL I&C) daughterboard for interfacing with the VME bus. and finally, the card will be packaged as a standard-height, single-wide (6U4HP format) VME card.

First 12 Beam Position Monitors have been delivered. 2 of the viewers are also here, Master Machine assures us
that they will catch-up. There is no impact on the schedule other than my desire to get these behind us, girder
assemblies will not start for ~ 3 months. A clean hood will be added to Area 51 in the FEL to allow for dust free
assembly of the "odds & ends". Some of these are Michelle's Ultra-Viewers!, and the wiggler viewers.

James continued to spend most of the week finishing up on the LSS system upgrades.  The new box in Lab 2 has
been installed and tested; the LSS Master Chassis has been upgraded to include all six user labs; plans are in
progress for adding additional protective beam stops to lab 3 for multiple hutches.  The hutch in lab 6 is being
equipped with multiple magnetic switches to ensure all doors on the hutch are shut prior to beam operation in that
lab when in the hutch mode.

Aside from attending the Low Level RF workshop hosted by JLab this week, I met with folks in town from
Triumf. They are using a CAN bus driven system to control their power supplies. This "TRIPS" system is exactly
the model we hope to use to control the new commercial trim power supply racks. Both Kepco and Mideastern
Industries are interested in bidding on the racks. Further discussions are continuing with the folks who did the
project; Daryl Bishop and Don Dale.  The boxes for the new OBPM's are in. These are being painted and
assembled. Testing will take place next week using the drive laser.

Documentation of the Picomotor Controller Chassis is complete.  EECAD is currently working on the Picomotor
Relay Chassis and ECOs for the LSS Master Chassis and LSS Lab Interface Box.

Work on the Ion Pump Power Supply Chassis continues.  Documentation is nearly complete and most of the
parts have been received.

All of the parts for the User Camera Power Supply Chassis are in as soon as the front and back panels are
machined and the silk screen complete assembly can take place.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o  DULY Research was finally successful in getting their magnetic model of the small injector Dipole (GV) to
    achieve the required tolerances of 1 part in 10,000 for both field flatness and field integral.  Their
    trick was to add current to one side to achieve field uniformity and to have sliding magnetic shorts on the field
    clamp to achieve field integral uniformity.
o They are working on incorporating the findings of the magnetic model into a practical hardware design.
o They continue work on the details of the coil for the GU.
Optical Chicane Dipole (DW)
o The design package is going through some small alterations.
o The contract was signed with Innovation Services to review our coil specification and bring it up to date for
    epoxy formulation, curing, manufacturing and quality control.
Arc Dipoles (GY, GX, GQ)
o We got he samples of electrical grade steel for the pole tips of the GW and GY.  We were able to select
    material that has a 20-inch sweet zone that is uniform to within 0.0001 inch. Sixty 22 inch long sheets of
    this material were put on order.
o  AES has stopped design detailing of the 180 degree dipole (GY) pending the final magnetic model findings.
    They are adding the hole for the synchrotron light periscopes.
o  The magnetic analysis of the GY with path length correctors and now synchrotron light periscopes continued.
    We found that there was a misunderstanding between the energy of the beam and the momentum of the beam
    (Our momentum based specification is 210 MeV/c) that resulted in Tom assuming a more energetic beam than
    our assumption.  In his particle tracking, this falsely overstated the beam offsets quoted in past reports.  At
    week’s end Tom Schultheiss is calculating parallel offset if a beam of less than 3.8 mm and is homing in on the
    proper face extension and Purcell Gap decrease values.
o They have begun design and magnetic modeling the Arc Bend Dipole (GX) after some go-rounds on the exact
QG (3 inch quad)
The coil-manufacturing package remains in Procurement cycle.
QX (Trim Quad)
o  We got a working magnetic model and are now debugging its meshing.
o  A vendor has agreed to custom make conductor for these quadrupoles to a defined dimension to a very high
    tolerance.  This should considerably decrease the manufacturing problems associated with conductor size that
    have plagued us in the past.
Corrector Dipoles (DB & DJ)
o We discovered that the drawings on the previous bid were sent out without a specification.   At week’s end a
    new package, (that will be easier for vendor’s to understand) is in vendor’s hands for a re-bid.
o The Dipole Steel is in Procurement cycle.
o We started design layout of the girders between cryomodule regions.
o We made considerable advances in firming up the parameters associated with dipole chamber height and the
    configurations of the arc chambers.
o We sent out inquires to vacuum vendors to determine if they can make our style of chamber.
o We firmed up our overall approach to piping the LCW.
o The alignment group at CERN is going to sell us 6 of their heavy weight rated cartridges (for a good price)
    for use under our 180° dipoles so we will not have to go through the engineering.

WBS 10 (Wiggler):
The dispersion section magnet was disassembled to add holes that the vendor missed onto the support stand.
The magnet was reassembled and setup on the magnet measurement stand.  Analysis of initial measurement
data in general shows a very high quality field.  It has a DC bias, similar to the one shown on the wiggler but of a
much smaller magnitude.

We will be exploring ways to short out or cancel this field in the coming week.  A coil design for the wiggler end
correctors has been submitted to the shop.  In the meantime we will be taking measurements of the DC field in
the wiggler vs. the strength of the end correction using some AT coils left over from CEBAF.  This should allow
us to determine the excitation necessary for these coils.  All components for the wiggler viewers have now been
received.  The small tubing for the vacuum chamber has also been received.  We should be able to start
machining of the chamber soon.

WBS 11 (Optics):
We had Robert Chow (LLNL) here on Tuesday to evaluate our plans for the Mirror Test Stand (MTS).
Overall, he thought we had a very good plan.  He did alert us to some experimental subtleties, and offered to
share some designs that proved successful in the high average power laser system (AVLIS) operated at LLNL in
the '80s & '90s.  We anticipate receipt of the rest of the MTS internal hardware next week.  We reviewed the
check prints for the new ultraviewer, and identified several changes; these are being incorporated.  We should
proceed to signoff next week.  Conceptual designs for the diagnostic beam dump and the collimator proceeds,
for the latter, S. Benson has done some optical modeling that defines the types of mirrors we'll need.

Parts for the manually-actuated turning mirror (initially for the User Lab 3) were received from our machine shop.

Assembling them identified a few minor interferences, these are being fixed.  Supports for the beamline are being
designed and procured.  Some work was done on the drive laser to realign it's output; it will be needed next
week for OBPM tests. The LSOP for User Lab 2 has been signed off.  After we receive some beamline shields,
we'll begin testing the laser, and then begin our first checkouts of the MTS.


No operations this week. Planning for the summer FEL user run continues.

Privacy and Security Notice